OpenStudy (anonymous):

Which set of rectangular coordinates describes the same location as the polar coordinates (4,3pi/2)?

OpenStudy (anonymous):

\[\begin{cases}x=r\cos\theta\\ y=r\sin\theta\end{cases}~~\iff~~\begin{cases}x=4\cos\dfrac{3\pi}{2}\\\\ y=4\sin\dfrac{3\pi}{2}\end{cases}~~\iff~~\begin{cases}x=0\\\\ y=-4\end{cases}\]